Consider the initial value problem y +3y +2y = (t-1)+r(t), y(0) = y(0) = 0, where 8(t-1) is Diracs delta function and S4 i
Answer #1

Page Nord r(t) = 4+ if ost<12 if t> 8 ret) H 4 t r(t) = 4+ ult) – 4 (4 -2) u (1-1) +44 (+-2) HitHH 1 4 4 u(t)-4 (1-2) u (t-1)Page No.2 LT of St-1)+r(t) 5(+-1) + 4+u(t) - 4 (4-2) u(t-1)+4 (t-1) Laplace Transform LT = € S + 4 4 es $2 + 4 S 32 @y+3y + 2Page Nory y(s) - -s es + 5 2 8 1 St ? + 4 St1 St2 52 st2 [2es -5es + 4ers les S 52 sti st2 + 2 e-s thens St1 S42 Ihverse L. T